Find the average of magnitude of linear momentum of helium molecules in sample of helium gas at temperature 150pi K .mass of helium molecules =166/3×10^-27kgand R=25/3j-mol k
1
Expert's answer
2019-03-07T08:37:41-0500

The average of magnitude of linear momentum of helium molecules


"p=mv_{\rm{ave}}"

The average speed of molecules


"v_{\rm{ave}}=\sqrt{\frac{8RT}{\pi M}}=\sqrt{\frac{8RT}{\pi mN_A}}"

So,


"p=m\sqrt{\frac{8RT}{\pi mN_A}}=\sqrt{\frac{8mRT}{\pi N_A}}"

"=\sqrt{\frac{8\times 166/3\times 10^{-27}\times 25/3\times 150\pi}{6\pi \times 10^{23}}}"

"=3\times 10^{-23}\;\rm{kg\cdot m/s}"